OpenStudy (katherinesmith):

Write the equation of a parabola with a vertex at (-3, 7) and a directrix x = -8.

OpenStudy (katherinesmith):

@tkhunny

OpenStudy (tkhunny):

With the vertex and directrix, the focus should be trivial. Where is it?

OpenStudy (tkhunny):

With the vertex and directrix, "p" should be trivial along with 4p. What is "p"?

OpenStudy (katherinesmith):

wait jump ahead a couple of steps. is the equation \[y - 7 = \frac{ 1 }{ 20 }(x + 3)^{2}\]

OpenStudy (tkhunny):

No. That opens UP and you should be opening RIGHT. You do have p and the vertex.

OpenStudy (katherinesmith):

so \[x + 3 = \frac{ 1 }{ 20 }(y - 7)^{2}\]

OpenStudy (tkhunny):

That's better. Do you believe it?

OpenStudy (katherinesmith):

i think so.

OpenStudy (tkhunny):

You can ALWAYS check with the definition. With directrix at x = -8 and focus at (2,7), an application of the definition of a parabola should have the same result. Good work.